OpenStudy (anonymous):

The product of two consecutive negative integers is 600. What is the value of the lesser integer? –60 –30 –25 –15

OpenStudy (alexandervonhumboldt2):

a=b+1 b*a=600

OpenStudy (alexandervonhumboldt2):

b*(b+1)=600

OpenStudy (anonymous):

thats the equation ?

OpenStudy (alexandervonhumboldt2):

no let me solve it

OpenStudy (anonymous):

Ok

OpenStudy (alexandervonhumboldt2):

@texaschic101 can you help i got confused?

ganeshie8 (ganeshie8):

guess and check is easy here

ganeshie8 (ganeshie8):

since the nearest perfect square is 625, the required numbers must be around -25

ganeshie8 (ganeshie8):

there is only one number around -25 in your options so essentially you're done but just verify whether -25 and -24 multiply to 600 or not